# Instruction Tuning for Large Language Models: A Survey

### v5

@@ Abstract
This paper surveys research works in the quickly advancing field of instruction tuning (IT), a crucial technique to enhance the capabilities and controllability of large language models (LLMs). Instruction tuning refers to the process of further training LLMs on a dataset consisting of (instruction, output) pairs in a supervised fashion, which bridges the gap between the next-word prediction objective of LLMs and the users’ objective of having LLMs adhere to human instructions. In this work, we make a systematic review of the literature, including the general methodology of IT, the construction of IT datasets, the training of IT models, and applications to different modalities, domains and application, along with analysis on aspects that influence the outcome of IT (e.g., generation of instruction outputs, size of the instruction dataset, etc). We also review the potential pitfalls of IT along with criticism against it, along with efforts pointing out current deficiencies of existing strategies and suggest some avenues for fruitful research. ††footnotetext: Zhejiang University, Shannon.AI, Nanyang Technological University, Amazon ††footnotetext: Email: sy_zhang@zju.edu.cn ††footnotetext: Project page can be found at: https://github.com/xiaoya-li/Instruction-Tuning-Survey ††footnotetext: * The latest update was on March 12, 2024.
@@ 1 Introduction
The field of large language models (LLMs) has witnessed remarkable progress in recent years. LLMs such as GPT-3 (Brown et al., 2020b), PaLM (Chowdhery et al., 2022), and LLaMA (Touvron et al., 2023a) have demonstrated impressive capabilities across a wide range of natural language tasks (Zhao et al., 2021; Wang et al., 2022b, 2023a; Wan et al., 2023; Sun et al., 2023c; Wei et al., 2023a; Li et al., 2023a; Gao et al., 2023a; Yao et al., 2023; Yang et al., 2022a; Qian et al., 2022; Lee et al., 2022; Yang et al., 2022b; Gao et al., 2023b; Ning et al., 2023; Liu et al., 2021b; Wiegreffe et al., 2021; Sun et al., 2023b, a; Adlakha et al., 2023; Chen et al., 2023b). One of the major issues with LLMs is the mismatch between the training objective and users’ objective: LLMs are typically trained on minimizing the contextual word prediction error on large corpora; while users want the model to "follow their instructions helpfully and safely" (Radford et al., 2019; Brown et al., 2020a; Fedus et al., 2021; Rae et al., 2021; Thoppilan et al., 2022)
@@
To address this mismatch, instruction tuning (IT) is proposed, serving as an effective technique to enhance the capabilities and controllability of large language models. It involves further training LLMs using (instruction, output) pairs, where instruction denotes the human instruction for the model, and output denotes the desired output that follows the instruction. The benefits of IT are threefold: (1) Finetuning an LLM on the instruction dataset bridges the gap between the next-word prediction objective of LLMs and the users’ objective of instruction following; (2) IT allows for a more controllable and predictable model behavior compared to standard LLMs. The instructions serve to constrain the model’s outputs to align with the desired response characteristics or domain knowledge, providing a channel for humans to intervene with the model’s behaviors; and (3) IT is computationally efficient and can help LLMs rapidly adapt to a specific domain without extensive retraining or architectural changes.
@@
Despite its effectiveness, IT also poses challenges: (1) Crafting high-quality instructions that properly cover the desired target behaviors is non-trivial: existing instruction datasets are usually limited in quantity, diversity, and creativity; (2) there has been an increasing concern that IT only improves on tasks that are heavily supported in the IT training dataset Gudibande et al. (2023); and (3) there has been an intense criticism that IT only captures surface-level patterns and styles (e.g., the output format) rather than comprehending and learning the task Kung and Peng (2023). Improving instruction adherence and handling unanticipated model responses remain open research problems. These challenges highlight the importance of further investigations, analysis, and summarization in this field, to optimize the fine-tuning process and better understand the behavior of instruction fine-tuned LLMs.
@@
In the literature, there has been an increasing research interest in analysis and discussions on LLMs, including pre-training methods Zhao et al. (2023), reasoning abilities Huang and Chang (2022), downstream applications Yang et al. (2023a); Sun et al. (2023b), but rarely on the topic of LLM instruction finetuning. This survey attempts to fill this blank, organizing the most up-to-date state of knowledge on this quickly advancing field. Specifically,
Section 2 presents the general methodology employed in instruction fine-tuning.
@@
Section 4 presents representative instruction-finetuned models.
@@
Section 7 reviews explorations to make instruction fine-tuning more efficient, reducing the computational and time costs associated with adapting large models.
@@ 2.1 Instruction Dataset Construction
2.2 Instruction Tuning
@@ 3.1.1 Natural Instructions
P3 (Public Pool of Prompts) (Sanh et al., 2021) is an instruction fine-tuning dataset constructed by integrating 170 English NLP datasets and 2,052 English prompts. Prompts, which are sometimes named task templates, are functions that map a data instance in a conventional NLP task (e.g., question answering, text classification) to a natural language input-output pair.
@@ 3.1.5 LIMA
LIMA (Zhou et al., 2023) is an English instruction dataset consisting of a train set with 1K data instances and a test set with 300 instances. The train set contains 1K ("instruction", "response") pairs. For the training data, 75% are sampled from three community question & answers websites (i.e., Stack Exchange, wikiHow, and the Pushshift Reddit Dataset (Baumgartner et al., 2020)); 20% are manually written by a set of the authors (referred Group A) inspired by their interests; 5% are sampled from the Super-Natural Instructions dataset (Wang et al., 2022d). As for the valid set, the authors sampled 50 instances from the Group A author-written set. The test set contains 300 examples, with 76.7% written by another group (Group B) of authors and 23.3% sampled from the Pushshift Reddit Dataset (Baumgartner et al., 2020), which is a collection of questions & answers within the Reddit community.
@@ 3.2 Synthetic Data via Distillation
In addition to the above datasets, there are many datasets in general domain, such as: ShareGPT111https://huggingface.co/datasets/RyokoAI/ShareGPT52K, WildChat Zhao et al. (2024), Vicuna Zheng et al. (2024), Unnatural Instructions (Honovich et al., 2022). Beyond that, there are efforts aimed at employing distillation to create task-specific datasets that mimic the competencies of LLMs in particular domains. For example, for coding generation, there are WizardCoder Luo et al. (2023), Magicoder Wei et al. (2023b) and WaveCoder Yu et al. (2023), for reasoning and writing, there are Phi-1 Gunasekar et al. (2023) and Phi-1.5 Li et al. (2023h), and for ranking, there is Nectar Zhu et al. (2023a).
@@ 3.3.2 Instruction Back-translation
4 Instruction Fine-tuned LLMs
@@
https://huggingface.co/bigscience/bloomz
@@
https://huggingface.co/google/flan-t5-xxl
@@
https://github.com/tatsu-lab/stanford_alpaca
@@
https://github.com/lm-sys/FastChat
@@
https://github.com/Instruction-Tuning-with-GPT-4/GPT-4-LLM
@@
https://github.com/nlpxucan/WizardLM
@@
https://github.com/THUDM/ChatGLM2-6B
@@
https://huggingface.co/facebook/opt-iml-30b
@@
https://github.com/databrickslabs/dolly
@@
https://huggingface.co/tiiuae/falcon-40b-instruct
@@
https://huggingface.co/JosephusCheung/Guanaco
@@
https://huggingface.co/openaccess-ai-collective/minotaur-15b
@@
https://huggingface.co/NousResearch/Nous-Hermes-13b
@@
https://github.com/allenai/open-instruct
@@
https://github.com/RUC-GSAI/YuLan-Chat
@@
https://github.com/OpenLMLab/MOSS
@@
https://github.com/jondurbin/airoboros
@@
https://github.com/thunlp/UltraChat
In this section, we detail widely-used LLM models in the community that are trained through instruction fine-tuning.
@@ 4.2 BLOOMZ
For automatic evaluation, BLOOMZ performs better than BLOOM in the zero-shot setting by +10.4%, 20.5%, and 9.8% on coreference resolution, sentence completion and natural language inference datasets, respectively. For the HumanEval benchmark (Chen et al., 2021), BLOOMZ outperforms BLOOM by 10% in terms of the Pass@100 metric. For generative tasks, BLOOMZ receives +9% BLEU improvement compared to BLOOM on the lm-evaluation-harness benchmark222https://github.com/EleutherAI/lm-evaluation-harness.
@@ 4.3 Flan-T5
For evaluation, FLAN-T5 (11B) outperforms T5 (11B), and achieves comparable results to larger models, including PaLM (60B) (Chowdhery et al., 2022) in the few-shot setting. FLAN-T5 outperforms T5 by +18.9%, +12.3%, +4.1%, +5.8%, +2.1%, and +8% on MMLU (Hendrycks et al., 2020), BBH (Suzgun et al., 2022), TyDiQA (Clark et al., 2020), MGSM (Shi et al., 2022), open-ended generation, and RealToxicityPrompts (Gehman et al., 2020), respectively. In few-shot settings, FLAN-T5 outperforms PaLM +1.4% and +1.2% on the BBH and TyDiQA datasets.
@@ 4.5 Vicuna
Vicuna (13B) (Chiang et al., 2023) is a language model trained by fine-tuning LLaMA (13B) (Touvron et al., 2023a) on the conversational dataset generated by ChatGPT333https://openai.com/blog/chatgpt.
The authors gathered user-shared ChatGPT conversations from ShareGPT.com444https://sharegpt.com/, and got 70K conversation records after filtering out low-quality samples. LLaMA (13B) was fine-tuned on the constructed conversation dataset using a modified loss function tailored to multi-turn conversations. To better understand long context across multiple-turn dialog, the authors expanded the max context length from 512 to 2048. For training, the authors adopted the gradient checkpointing and flash attention (Dao et al., 2022) techniques to reduce the GPU memory cost in the fine-tuning process. The fine-tuning process takes 24 hours on an 8 80GB A100 device with fully shared data parallelism.
@@ 4.7 Claude
Claude555https://www.anthropic.com/index/introducing-claude is a language model trained by fine-tuning the pre-trained language model on an instruction dataset, aiming to generate helpful and harmless responses. The fine-tuning process consists of two stages: (1) supervised fine-tuning on the instruction dataset. The authors created an instruction dataset by collecting 52K different instructions, paired with responses generated by GPT-4. The fine-tuning process takes approximately eight hours on an 8-card 80GB A100 machine with mixed precision and fully shared data parallelism. (2) optimizing the step-1 model with the proximal policy optimization (Schulman et al., 2017) method. The authors first built a comparison dataset by collecting responses from multiple large language models (e.g., GPT-3 (Brown et al., 2020b)) to the given collection of instructions and then asking GPT-4 (OpenAI, 2023) to rate each response. Using the ratings, a reward model is trained. Then, the fine-tuned model from Step 1 is optimized using the reward model with the proximal policy optimization method.
@@ 4.9 ChatGLM2
Evaluations are conducted on four English and Chinese benchmarks, including MMLU (English) (Hendrycks et al., 2020), C-Eval (Chinese) (Huang et al., 2023), GSM8K (Math) (Cobbe et al., 2021), and BBH (English) (Suzgun et al., 2022). ChatGLM2 (6B) outperforms GLM (6B) and the baseline model ChatGLM (6B) on all benchmarks. Specifically, ChatGLM2 outperforms GLM by +3.1 on MMLU, +5.0 on C-Eval, +8.6 on GSM8K, and +2.2 on BBH. ChatGLM2 achieves better performances than ChatGLM by +2.1, +1.2, +0.4, +0.8 on MMLU, C-Eval, GSM8K and BBH, respectively.
@@ 4.10 LIMA
LIMA (65B) (Zhou et al., 2023) is a large language model trained by fine-tuning LLaMA (65B) (Touvron et al., 2023a) on an instruction dataset, which is constructed based on the proposed superficial alignment hypothesis.
The superficial alignment hypothesis refers to the idea that the knowledge and capabilities of a model are almost acquired in the pre-training stage, while the alignment training (e.g., instruction fine-tuning) teaches models to generate responses under user-preferred formalizations. Based on the superficial alignment hypothesis, the authors claimed that large language models can generate user-satisfied responses by fine-tuning it on a small fraction of instruction data. Therefore, the authors built instruction train/valid/test sets to verify this hypothesis.
Evaluations are conducted on the constructed test set. For human evaluations, LIMA outperforms InstructGPT and Alpaca by 17% and 19%, respectively. Additionally, LIMA achieves comparable results to BARD666https://bard.google.com/, Cladue777https://www.anthropic.com/index/introducing-claude, and GPT-4. For automatic evaluation, which is conducted by asking GPT-4 to rate responses and a higher rate score denotes better performance, LIMA outperforms InstructGPT and Alpaca by 20% and 36%, respectively, achieving comparable results to BARD, while underperforming Claude and GPT-4. Experimental results verify the proposed superficial alignment hypothesis.

### v7

@@ Abstract
This paper surveys research works in the quickly advancing field of instruction tuning (IT), which can also be referred to as supervised fine-tuning (SFT)111In this paper, unless specified otherwise, instruction tuning (IT) will be equivalent to supervised fine-tuning (SFT)., a crucial technique to enhance the capabilities and controllability of large language models (LLMs). Instruction tuning refers to the process of further training LLMs on a dataset consisting of (instruction, output) pairs in a supervised fashion, which bridges the gap between the next-word prediction objective of LLMs and the users’ objective of having LLMs adhere to human instructions. In this work, we make a systematic review of the literature, including the general methodology of IT, the construction of IT datasets, the training of IT models, and applications to different modalities, domains and application, along with analysis on aspects that influence the outcome of IT (e.g., generation of instruction outputs, size of the instruction dataset, etc). We also review the potential pitfalls of IT along with criticism against it, along with efforts pointing out current deficiencies of existing strategies and suggest some avenues for fruitful research. ††footnotetext: ♠Zhejiang University, ♣Shannon.AI, ▲Nanyang Technological University, ◆Amazon ††footnotetext: Email: sy_zhang@zju.edu.cn ††footnotetext: Project page can be found at: https://github.com/xiaoya-li/Instruction-Tuning-Survey ††footnotetext: * The latest update was on October 16, 2024.
@@ 1 Introduction
The field of large language models (LLMs) has witnessed remarkable progress in recent years. LLMs such as GPT-3 (Brown et al., 2020b), PaLM (Chowdhery et al., 2022), and LLaMA (Touvron et al., 2023a) have demonstrated impressive capabilities across a wide range of natural language tasks (Zhao et al., 2021; Wang et al., 2022b, 2023b; Wan et al., 2023; Sun et al., 2023c; Wei et al., 2023a; Li et al., 2023a; Gao et al., 2023a; Yao et al., 2023; Yang et al., 2022a; Qian et al., 2022; Lee et al., 2022; Yang et al., 2022b; Gao et al., 2023b; Ning et al., 2023; Liu et al., 2021b; Wiegreffe et al., 2021; Sun et al., 2023b, a; Adlakha et al., 2023; Chen et al., 2023b). One of the major issues with LLMs is the mismatch between the training objective and users’ objective: LLMs are typically trained on minimizing the contextual word prediction error on large corpora; while users want the model to "follow their instructions helpfully and safely" (Radford et al., 2019; Brown et al., 2020a; Fedus et al., 2021; Rae et al., 2021; Thoppilan et al., 2022)
@@
To address this mismatch, instruction tuning (IT), which can also be referred to as supervised fine-tuning (SFT), is proposed, serving as an effective technique to enhance the capabilities and controllability of large language models. It involves further training LLMs using (instruction, output) pairs, where instruction denotes the human instruction for the model, and output denotes the desired output that follows the instruction. The benefits of IT are threefold: (1) Finetuning an LLM on the instruction dataset bridges the gap between the next-word prediction objective of LLMs and the users’ objective of instruction following; (2) IT allows for a more controllable and predictable model behavior compared to standard LLMs. The instructions serve to constrain the model’s outputs to align with the desired response characteristics or domain knowledge, providing a channel for humans to intervene with the model’s behaviors; and (3) IT is computationally efficient and can help LLMs rapidly adapt to a specific domain without extensive retraining or architectural changes.
@@
Despite its effectiveness, IT also poses challenges: (1) Crafting high-quality instructions that properly cover the desired target behaviors is non-trivial: existing instruction datasets are usually limited in quantity, diversity, and creativity; (2) there has been an increasing concern that IT only improves on tasks that are heavily supported in the IT training dataset Gudibande et al. (2023); and (3) there has been an intense criticism that IT only captures surface-level patterns and styles (e.g., the output format) rather than comprehending and learning the task Kung and Peng (2023). Improving instruction adherence and handling unanticipated model responses remain open research problems. These challenges highlight the importance of further investigations, analysis, and summarization in this field, to optimize the fine-tuning process and better understand the behavior of instruction tuned LLMs.
@@
In the literature, there has been an increasing research interest in analysis and discussions on LLMs, including pre-training methods Zhao et al. (2023), reasoning abilities Huang and Chang (2022), downstream applications Yang et al. (2023a); Sun et al. (2023b), but rarely on the topic of LLM instruction tuning. This survey attempts to fill this blank, organizing the most up-to-date state of knowledge on this quickly advancing field. Specifically,
Section 2 presents the general methodology employed in instruction tuning.
@@
Section 4 presents representative instruction tuned models.
@@
Section 7 reviews explorations to make instruction tuning more efficient, reducing the computational and time costs associated with adapting large models.
@@ 2.1 Instruction Dataset Construction
2.2 Instruction Tuning / Supervised Fine-tuning
@@ 3.1.1 Natural Instructions
P3 (Public Pool of Prompts) (Sanh et al., 2021) is an instruction tuning dataset constructed by integrating 170 English NLP datasets and 2,052 English prompts. Prompts, which are sometimes named task templates, are functions that map a data instance in a conventional NLP task (e.g., question answering, text classification) to a natural language input-output pair.
@@ 3.1.5 LIMA
LIMA (Zhou et al., 2023a) is an English instruction dataset consisting of a train set with 1K data instances and a test set with 300 instances. The train set contains 1K ("instruction", "response") pairs. For the training data, 75% are sampled from three community question & answers websites (i.e., Stack Exchange, wikiHow, and the Pushshift Reddit Dataset (Baumgartner et al., 2020)); 20% are manually written by a set of the authors (referred Group A) inspired by their interests; 5% are sampled from the Super-Natural Instructions dataset (Wang et al., 2022d). As for the valid set, the authors sampled 50 instances from the Group A author-written set. The test set contains 300 examples, with 76.7% written by another group (Group B) of authors and 23.3% sampled from the Pushshift Reddit Dataset (Baumgartner et al., 2020), which is a collection of questions & answers within the Reddit community.
@@ 3.2 Synthetic Data via Distillation
In addition to the above datasets, there are many datasets in general domain, such as: ShareGPT222https://huggingface.co/datasets/RyokoAI/ShareGPT52K, WildChat Zhao et al. (2024), Vicuna Zheng et al. (2024), Unnatural Instructions (Honovich et al., 2022). Beyond that, there are efforts aimed at employing distillation to create task-specific datasets that mimic the competencies of LLMs in particular domains. For example, for coding generation, there are WizardCoder Luo et al. (2023), Magicoder Wei et al. (2023b) and WaveCoder Yu et al. (2023), for reasoning and writing, there are Phi-1 Gunasekar et al. (2023) and Phi-1.5 Li et al. (2023i), and for ranking, there is Nectar Zhu et al. (2023a).
@@ 3.3.2 Instruction Back-translation
4 Instruction Tuned LLMs
@@
1 https://huggingface.co/bigscience/bloomz
@@
2 https://huggingface.co/google/flan-t5-xxl
@@
3 https://github.com/tatsu-lab/stanford_alpaca
@@
4 https://github.com/lm-sys/FastChat
@@
5 https://github.com/Instruction-Tuning-with-GPT-4/GPT-4-LLM
@@
6 https://github.com/nlpxucan/WizardLM
@@
7 https://github.com/THUDM/ChatGLM2-6B
@@
8 https://huggingface.co/facebook/opt-iml-30b
@@
9 https://github.com/databrickslabs/dolly
@@
10 https://huggingface.co/tiiuae/falcon-40b-instruct
@@
11 https://huggingface.co/JosephusCheung/Guanaco
@@
12 https://huggingface.co/openaccess-ai-collective/minotaur-15b
@@
13 https://huggingface.co/NousResearch/Nous-Hermes-13b
@@
14 https://github.com/allenai/open-instruct
@@
15 https://github.com/RUC-GSAI/YuLan-Chat
@@
16 https://github.com/OpenLMLab/MOSS
@@
17 https://github.com/jondurbin/airoboros
@@
18 https://github.com/thunlp/UltraChat
In this section, we detail widely-used LLM models in the community that are trained through instruction tuning.
@@ 4.2 BLOOMZ
For automatic evaluation, BLOOMZ performs better than BLOOM in the zero-shot setting by +10.4%, 20.5%, and 9.8% on coreference resolution, sentence completion and natural language inference datasets, respectively. For the HumanEval benchmark (Chen et al., 2021b), BLOOMZ outperforms BLOOM by 10% in terms of the Pass@100 metric. For generative tasks, BLOOMZ receives +9% BLEU improvement compared to BLOOM on the lm-evaluation-harness benchmark333https://github.com/EleutherAI/lm-evaluation-harness.
@@ 4.3 Flan-T5
For evaluation, FLAN-T5 (11B) outperforms T5 (11B), and achieves comparable results to larger models, including PaLM (60B) (Chowdhery et al., 2022) in the few-shot setting. FLAN-T5 outperforms T5 by +18.9%, +12.3%, +4.1%, +5.8%, +2.1%, and +8% on MMLU (Hendrycks et al., 2020b), BBH (Suzgun et al., 2022b), TyDiQA (Clark et al., 2020), MGSM (Shi et al., 2022), open-ended generation, and RealToxicityPrompts (Gehman et al., 2020), respectively. In few-shot settings, FLAN-T5 outperforms PaLM +1.4% and +1.2% on the BBH and TyDiQA datasets.
@@ 4.5 Vicuna
Vicuna (13B) (Chiang et al., 2023) is a language model trained by fine-tuning LLaMA (13B) (Touvron et al., 2023a) on the conversational dataset generated by ChatGPT444https://openai.com/blog/chatgpt.
The authors gathered user-shared ChatGPT conversations from ShareGPT.com555https://sharegpt.com/, and got 70K conversation records after filtering out low-quality samples. LLaMA (13B) was fine-tuned on the constructed conversation dataset using a modified loss function tailored to multi-turn conversations. To better understand long context across multiple-turn dialog, the authors expanded the max context length from 512 to 2048. For training, the authors adopted the gradient checkpointing and flash attention (Dao et al., 2022) techniques to reduce the GPU memory cost in the fine-tuning process. The fine-tuning process takes 24 hours on an 8 80GB A100 device with fully shared data parallelism.
@@ 4.7 Claude
Claude666https://www.anthropic.com/index/introducing-claude is a language model trained by fine-tuning the pre-trained language model on an instruction dataset, aiming to generate helpful and harmless responses. The fine-tuning process consists of two stages: (1) supervised fine-tuning on the instruction dataset. The authors created an instruction dataset by collecting 52K different instructions, paired with responses generated by GPT-4. The fine-tuning process takes approximately eight hours on an 8-card 80GB A100 machine with mixed precision and fully shared data parallelism. (2) optimizing the step-1 model with the proximal policy optimization (Schulman et al., 2017) method. The authors first built a comparison dataset by collecting responses from multiple large language models (e.g., GPT-3 (Brown et al., 2020b)) to the given collection of instructions and then asking GPT-4 (OpenAI, 2023) to rate each response. Using the ratings, a reward model is trained. Then, the fine-tuned model from Step 1 is optimized using the reward model with the proximal policy optimization method.
@@ 4.9 ChatGLM2
Evaluations are conducted on four English and Chinese benchmarks, including MMLU (English) (Hendrycks et al., 2020b), C-Eval (Chinese) (Huang et al., 2023), GSM8K (Math) (Cobbe et al., 2021), and BBH (English) (Suzgun et al., 2022b). ChatGLM2 (6B) outperforms GLM (6B) and the baseline model ChatGLM (6B) on all benchmarks. Specifically, ChatGLM2 outperforms GLM by +3.1 on MMLU, +5.0 on C-Eval, +8.6 on GSM8K, and +2.2 on BBH. ChatGLM2 achieves better performances than ChatGLM by +2.1, +1.2, +0.4, +0.8 on MMLU, C-Eval, GSM8K and BBH, respectively.
@@ 4.10 LIMA
LIMA (65B) (Zhou et al., 2023a) is a large language model trained by fine-tuning LLaMA (65B) (Touvron et al., 2023a) on an instruction dataset, which is constructed based on the proposed superficial alignment hypothesis.
The superficial alignment hypothesis refers to the idea that the knowledge and capabilities of a model are almost acquired in the pre-training stage, while the alignment training (e.g., instruction tuning) teaches models to generate responses under user-preferred formalizations. Based on the superficial alignment hypothesis, the authors claimed that large language models can generate user-satisfied responses by fine-tuning it on a small fraction of instruction data. Therefore, the authors built instruction train/valid/test sets to verify this hypothesis.
Evaluations are conducted on the constructed test set. For human evaluations, LIMA outperforms InstructGPT and Alpaca by 17% and 19%, respectively. Additionally, LIMA achieves comparable results to BARD777https://bard.google.com/, Cladue888https://www.anthropic.com/index/introducing-claude, and GPT-4. For automatic evaluation, which is conducted by asking GPT-4 to rate responses and a higher rate score denotes better performance, LIMA outperforms InstructGPT and Alpaca by 20% and 36%, respectively, achieving comparable results to BARD, while underperforming Claude and GPT-4. Experimental results verify the proposed superficial alignment hypothesis.